How many moles of Hg are contained in 100ml of Hg?
1
Expert's answer
2017-04-13T10:51:25-0400
Number of moles (n) is equal to the mass of the sample (m) divided by the molar mass of the substance (M):
n=m / M.

Mass of the sample (m) is equal to the volume of the sample (V) multiplied by density of the substance (D):
m = V * D.

So, finally we have:
n=(V * D) / M.

Molar mass of Hg is 200.6 g/mol, density of Hg is 13.5 g/cm3
https://en.wikipedia.org/wiki/Mercury_(element)
Note that 1 ml = 1 cm3.

Do the calculation:
n=(100 cm3 * 13.5 g/cm3) / 200.6 g/mol = 6.7 mol.


Answer:
6.7 moles of mercury.